OpenStudy (anonymous):

evaluate the expressin when x=2 1/3 and y= 3 1/5 . y=2 3/5

Directrix (directrix):

Is x=2 1/3 and y= 3 1/5 ...... x = (21)/3 or 7/3 and y = (31/5) or 16/5 ? Please clarify.

Directrix (directrix):

@ princessw --> Need your input